# ReCoRe: Regularized Contrastive Representation Learning of World Model

### v1

@@ Abstract
While recent model-free Reinforcement Learning (RL) methods have demonstrated human-level effectiveness in gaming environments, their success in everyday tasks like visual navigation has been limited, particularly under significant appearance variations. This limitation arises from (i) poor sample efficiency and (ii) over-fitting to training scenarios. To address these challenges, we present a world model that learns invariant features using (i) contrastive unsupervised learning and (ii) an intervention-invariant regularizer. Learning an explicit representation of the world dynamics i.e. a world model, improves sample efficiency while contrastive learning implicitly enforces learning of invariant features, which improves generalization. However, the naïve integration of contrastive loss to world models fails due to a lack of supervisory signals to the visual encoder, as world-model-based RL methods independently optimize representation learning and agent policy. To overcome this issue, we propose an intervention-invariant regularizer in the form of an auxiliary task such as depth prediction, image denoising, etc., that explicitly enforces invariance to style-interventions. Our method outperforms current state-of-the-art model-based and model-free RL methods and significantly on out-of-distribution point navigation task evaluated on the iGibson benchmark. We further demonstrate that our approach, with only visual observations, outperforms recent language-guided foundation models for point navigation, which is essential for deployment on robots with limited computation capabilities. Finally, we demonstrate that our proposed model excels at the sim-to-real transfer of its perception module on Gibson benchmark.
@@ 1 Introduction
In recent years, deep RL algorithms have been successfully employed for designing optimal strategies for games [25, 31] and shown a promise for controlling robots [21, 35]. Model-free RL approaches learn the policy along with the visual encoder in an end-to-end fashion from raw observations. As a result they require a large number of training samples, which makes it difficult to deploy them on real robots where obtaining a large amount of training data is resource intensive, especially for safety critical tasks such as autonomous navigation [9]. On the contrary, in model-based RL, an explicit predictive model of the world is learned i.e. world model, enabling the agent to plan by thinking ahead [6, 31, 10, 13]. The world model is learned separately from the policy, therefore, the policy can use the world model as a surrogate for real world. Consequently, model-based methods have higher sample efficiency [10, 13] making them more suitable in real environments since they can be trained with a small amount of data.
@@
Figure 1: Out-of-distribution depth prediction from world model. ReCoRe learns robust representations that are invariant to appearance variations which help in generalization of downstream task. Notice the consistent depth predictions despite texture variations on iGibson evaluation set.
@@
Nevertheless, even current model-based RL struggles with generalization, as model-based approaches have to learn the world model purely from experience, which poses several challenges. The central issue is the training bias in the model, which can be exploited by an agent, which leads to poor performance when deployed [10]. Another issue is that the latent representation is learned from a reconstruction loss, for example, the state abstraction of variational autoencoders (VAE) [18], is not sufficient to separate the task relevant states from irrelevant ones. Hence, the RL policy may still over-fit to environment-specific characteristics [40]. Thus, the aim of this paper is invariant features abstraction, if not the truly causal features, which is essential for learning robust RL policy.
@@
Figure 2: Flow diagram of proposed Regularized Contrastive Representation learning (ReCoRe) of World Model. It consists of four components: (i) invariant representation learning module, (ii) intervention-invariant regularization, and (iii) latent dynamics model and (iv) actor-critic controller. Invariant representation learning module utilizes data augmentation and contrastive learning for invariant features ($\tilde{s}_{t}$) abstraction from image observations ($o_{t}$). Latent dynamics model employs a recurrent neural network with deterministic hidden states ($h_{t}$) to predict the stochastic latent prior states ($\widehat{s}_{t}$), and corresponding rewards ($\widehat{r}_{t}$) from the posterior ($s_{t}$). Intervention invariant regularizer considers an auxilary task (here depth prediction i.e. $p_{\theta}(\widehat{d}_{t}|s_{t},h_{t})$ invariant to data augmentation (here texture variations) which prevents feature collapse in training world-model with contrastive learning. The controller maximizes the expected rewards of the action using an actor critic approach. Since the controller is learnt separately using invariant states of the environment, the proposed approach is more robust to out-of-distribution and sim-to-real generalization in addition to being sample efficient.
@@
In this paper, we propose to use contrastive learning for invariant state abstraction since the contrastive learning objective explicitly ensures that the feature embeddings are invariant to the intervention, i.e. data augmentation. However, feature collapse is possible if a naïve implementation is used as it can be seen from Table 1, that naïve implementation of contrastive loss (ReCoRe-D) completely fails ($<$1% success rate) for model-based RL. Mitrovic et al. [24] proposed an implicit regularizer based on KL-divergence that matches the distributions among the augmentations, which stabilises contrastive learning under model-free RL settings and slightly improved the performance over CURL [20]. On the contrary, we propose a regularizer in the form of an auxiliary task to explicitly enforce the invariant feature learning. Specifically, for navigation task we utilize depth predictions to extract the geometric features needed for navigation, reconstruction and other possible downstream tasks, as they do not dependent on textures as shown in Figure 1. We emphasize that depth is only required for training and not for deployment, since it works as a regulariser. Furthermore, in cases where depth is not available other auxilary tasks such as image denosing or optical flow prediction can be utilized for regularization, enabling a wider applicability of the proposed model. The design of interventions and invariant regularizers in the form of auxiliary losses becomes an interesting research problem which we will address in future work. Importantly, our setup allows us to use popular contrastive learning on model-based RL methods which improves the sample efficiency as well as helps in Out-of-Distribution (OoD) generalization.
@@
In summary, we propose a Regularized Contrastive Representation learning (ReCoRe) approach to the world model. The proposed variant of the world model can extract and predict the robust invariant features (Figure 2). ReCoRe is verified on the point goal navigation task from Gibson [37] and iGibson 1.0 [30] as well as the DeepMind control suite (DMControl) [33]. Our main contributions are:
@@
    We show that contrastive unsupervised representation learning can significantly improve OoD generalization of world model based reinforcement learning (refer Table 1, ReCoRe vs. DreamerV2).
@@
    We propose an intervention-invariant regularizer to learn the invariant features (Section 3.1.1) that is crucial in preventing feature collapse during training of world model based reinforcement learning approaches (as can be seen in Table 1, ReCoRe vs. ReCoRe-D).
@@
    Through extensive experiments, we showcase that our approach outperforms state-of-the-art RL models (including language guided foundation model Grounding DINO [23]) on out-of-distribution generalization (Table 1) and sim-to-real transfer of learned features (Table 2). We further show that even for in-distribution evaluation our approach is on par with model-free reinforcement learning approaches (Table 3) which is difficult for other model-based learning approaches.
@@ 2 Related Work
Unsupervised Representation Learning. Learning reusable feature representations from large unlabeled data has been an active research area. In the context of computer vision, one can leverage unlabeled images and videos to learn good intermediate representations, which can be useful for a wide variety of downstream tasks. Recently, VAE [18] has been a preferred approach for representation learning in model-based RL [10]. Since VAE does not make any additional consideration of downstream tasks, invariant representation learning with contrastive loss has shown more promising results [1, 20]. Self-supervised learning formulates representation learning as a supervised loss function between different transformations of data. In image-based learning self-supervision can be formulated using different image augmentations, for example, image distortion and rotation [7, 3]. We also use different data augmentation techniques to learn the invariant features using contrastive loss. Recently, transformer based visual models such as DINO [2] have shown to intrinsically capture robust object representations through self-supervision. Combining such models with language models pretrained on large corpus of data [22, 23] have resulted in extremely powerful representations that generalise well to diverse environments. However, inference on such large models is computationally expensive which makes it difficult to deploy on low budget robots. We also compare our proposed approach with Grounding DINO [23] features and showcase superior results, which we believe is the strongest baseline.
Contrastive Learning. Representation learning methods based on contrastive loss [4] have achieved state-of-the-art performance on face verification tasks. These methods use a contrastive loss to learn representations invariant to data augmentation [3, 16]. Given a list of input samples, contrastive loss forces samples from the same class to have similar embeddings and different ones for different classes. Since class labels are not available in the unsupervised setting, contrastive loss forces similar embedding for the augmented version of the same sample and different ones for different samples. There are several ways of formulating the contrastive loss such as Siamese [4], InfoNCE [34] and SimCLR [3]. We chose InfoNCE [34] for our contrastive loss in this work.
@@
Model-based RL. The human brain discovers the hidden causes underlying an observation. Those internal representations of the world influence how agents infer which actions will lead to a higher reward [15]. An early example of this idea was put forward by Sutton [32], where future hallucination samples rolled out from the learned world model are used in addition to the agent’s interactions for sample efficient learning. Further, planning through the world model has been successfully demonstrated in the world model by Ha et al. [10] and DreamerV2 by Hafner et al. [13]. Recently, replacing state extraction [29] and dynamic prediction [27] using transformer architecture is a popular direction, which further improves the results [29]. Masked World Models (MWM) combines the transformer based masked auto-encoder with DreamerV2. Other Dreamer variants DayDreamer [36] and DreamerV3 [14] aim to scale-up DreamerV2 architecture for physical robots and Atari games. DayDreamer [36] simply trains DreamerV2 on physical robots. While, DreamerV3 [14] proposes a few design choices such as symlog scaling, EMA regularization etc. that helps in scaling-up DreamerV2 to several domains. In our work, we propose to learn invariant features to improve OoD generalization and the sample efficiency further. While we utilize similar architecture as DreamerV2 for policy and world-model, our method is complementary to DreamerV3 and can take advantage from their design choices.
Sample Efficiency. Joint learning of auxiliary tasks with model-free RL makes them competitive with model-based RL in terms of sample efficiency. For example, the recently proposed model-free RL method called CURL [20] added contrastive loss as an auxiliary task and outperformed the state-of-the-art model-based RL method called Dreamer [12]. Also, two recent works using data augmentation for RL called RAD [19] and DrQ [38] outperform CURL without using an auxiliary contrastive loss. These results warrant that if an agent has access to a rich stream of data from the environments, an additional regularizer is unnecessary, since directly optimizing the policy objective is better than optimizing multiple objectives. However, we do not have access to a rich stream of data for many complex problems, hence sample efficiency still matters. Further, these papers do not consider the effect of regularizers in the form of auxiliary tasks and unsupervised representation learning for model-based RL, which is the main focus of our work.
@@ 3 ReCoRe-based World Model
We consider the visual control task as a finite-horizon partially observable Markov decision process (POMDP). We denote observation space, action space and time horizon as $\mathcal{O}$, $\mathcal{A}$ and $\mathcal{T}$ respectively. An agent performs continuous actions $a_{t}\sim p(a_{t}|o_{\leq t},a_{<t})$, and receives observations and scalar rewards $o_{t},r_{t}\sim p(o_{t},r_{t}|o_{<t},a_{<t})$ from the unknown environment. The goal of an agent is to maximize the expected total rewards $E_{p}(\sum_{t=1}^{T}r_{t})$. In the following sections we detail our proposed model.
@@ 3.1 World Model Design
We propose Regularized Contrastive Representation learning (ReCoRe) technique to learn the world model. Our method consists of four main components: (i) invariant representation learning module, (ii) intervention-invariant regularizer, (iii) latent dynamics model, and (iv) the controller. The data flow diagram of our proposed world model with explicitly regularized invariant feature learning is shown in Figure 2. We next describe the main components of our approach.
@@ 3.1.1 Invariant Representations Learning Module
Extracting representations that are invariant to appearance variations from image observations is a key component of our model. This, is crucial for improving robustness and out-of-distribution generalization of the model in real world. We learn these invariant representations by maximizing agreement between different style interventions of the same observation via a contrastive loss in the latent feature space. The world model optimizes feature learning and controller separately to improve the sample efficiency and simplify controller learning. We note that, most of the complexity of model-based RL approaches resides in the world model (i.e. the feature extraction and the dynamics model). Hence, we need a stronger supervisory signal to learn a better state representation of the environment. In this work, we used InfoNCE [34] style loss to learn invariant features. Hence, our encoder takes RGB observations ($o_{t}$), auxiliary inputs such as depth ($o_{t}^{d}$) while rewards ($r_{t}$) as inputs. The invariant feature learning is enforced by contrastive loss ($\mathcal{L}^{q}_{t}$). The proposed invariant features learning technique has following three sub-modules:
  • A style intervention module that utilizes data augmentation. We use spatial jitter, Gaussian blur, color jitter, grayscale and cutout data augmentation techniques for style intervention. Spatial jitter is implemented by first padding and then performing random crop. Given any observation $o_{t}$, our style intervention module randomly transforms it into two correlated views of the same observations. All the hyperparameters are provided in the appendix.
  • We use an encoder network that extracts representations from augmented observations, which is $\tilde{s_{t}}=encoder(o_{t})$. The encoder is optimized using contrastive loss (Equation 1), which by construction make the encoder invariant to the augmentations. We additionaly employ an EMA regularization [16] on the encoder for stabilizing training. The hyperparameters of the encoder is provided in the appendix.
  • Contrastive loss is defined for a contrastive prediction task, which can be explained as a differentiable dictionary lookup task. Given a query observation $q$ and a set of keys $K=\{k_{0},k_{1},...k_{2B}\}$ of length $2B$, with known positive $\{k_{+}\}$ and negative $\{k_{-}\}$ keys, the aim of contrastive loss is to learn a representation in which positive sample pairs stay close to each other while negative ones are far apart. In contrastive learning literature $q$, $K$, $k_{+}$ and ${k_{-}}$ are also referred as anchors, targets, positive and negative samples. We use bilinear products for projection head $W$ and InfoNCE loss for contrastive learning [34], which enforces the desired similarity in the embedding space:

### v2

@@ Abstract
While recent model-free Reinforcement Learning (RL) methods have demonstrated human-level effectiveness in gaming environments, their success in everyday tasks like visual navigation has been limited, particularly under significant appearance variations. This limitation arises from (i) poor sample efficiency and (ii) over-fitting to training scenarios. To address these challenges, we present a world model that learns invariant features using (i) contrastive unsupervised learning and (ii) an intervention-invariant regularizer. Learning an explicit representation of the world dynamics i.e. a world model, improves sample efficiency while contrastive learning implicitly enforces learning of invariant features, which improves generalization. However, the naïve integration of contrastive loss to world models is not good enough, as world-model-based RL methods independently optimize representation learning and agent policy. To overcome this issue, we propose an intervention-invariant regularizer in the form of an auxiliary task such as depth prediction, image denoising, image segmentation, etc., that explicitly enforces invariance to style interventions. Our method outperforms current state-of-the-art model-based and model-free RL methods and significantly improves on out-of-distribution point navigation tasks evaluated on the iGibson benchmark. With only visual observations, we further demonstrate that our approach outperforms recent language-guided foundation models for point navigation, which is essential for deployment on robots with limited computation capabilities. Finally, we demonstrate that our proposed model excels at the sim-to-real transfer of its perception module on the Gibson benchmark.
@@ 1 Introduction
In recent years, deep RL algorithms have been successfully employed for designing optimal strategies for games [25, 31] and shown a promise for controlling robots [21, 35]. Model-free RL approaches learn the policy along with the visual encoder in an end-to-end fashion from raw observations. As a result, they require a large number of training samples, which makes it difficult to deploy them on real robots where obtaining a large amount of training data is resource intensive, especially for safety-critical tasks such as autonomous navigation [9]. On the contrary, in model-based RL, an explicit predictive model of the world is learned called world model, enabling the agent to plan by thinking ahead [6, 31, 10, 13]. The world model is learned separately from the policy, therefore, the policy can use the world model as a surrogate for the real world. Consequently, model-based methods have higher sample efficiency [10, 13] making them more suitable in real environments since they can be trained with a small amount of data.
@@
Figure 1: Intervention invariant regularizer is applied in addition to the contrastive loss in world-model-based RL. ReCoRe learns robust representations that are invariant to out-of-distribution appearance variations which help in the generalization of downstream tasks (such as navigation). Notice the consistent depth predictions despite texture variations of an iGibson evaluation scene.
@@
Nevertheless, even current model-based RL struggles with generalization, as model-based approaches have to learn the world model purely from experience, which poses several challenges: The central issue is the training bias, which can be exploited by an agent, and leads to poor performance when deployed [10]. Another issue is that the latent representation is learned from a reconstruction loss, such as the state abstraction of variational autoencoders (VAE) [18], which is not sufficient to separate the task-relevant states from irrelevant ones. Hence, the RL policy may still overfit to environment-specific characteristics [40]. Thus, the aim of this paper is invariant feature abstraction, which is essential for learning a robust RL policy.
@@
Figure 2: Flow diagram of proposed Regularized Contrastive Representation learning (ReCoRe) of World Model. It consists of four components: (i) invariant representation learning module, (ii) intervention-invariant regularizer, (iii) latent dynamics model, and (iv) actor-critic controller. The invariant representation learning module utilizes data augmentation and contrastive learning for invariant features abstraction ($\tilde{s}_{t}$) from image observations ($o_{t}$). The latent dynamics model employs a recurrent neural network with deterministic hidden states ($h_{t}$) to predict the stochastic latent prior states ($\widehat{s}_{t}$), and corresponding rewards ($\widehat{r}_{t}$) from the posterior ($s_{t}$). Intervention invariant regularizer considers an auxiliary task (here depth prediction i.e. $p_{\theta}(\widehat{d}_{t}|s_{t},h_{t})$) invariant to data augmentation (here texture variations) which prevents feature collapse in training the world-model with contrastive learning. The controller maximizes the expected rewards of the action using an actor critic approach. In addition to being sample efficient, the proposed approach is more robust to out-of-distribution and sim-to-real generalization, since the controller is learned separately using invariant states of the environment.
@@
Specifically, we propose to use contrastive learning for invariant state abstraction since the contrastive learning objective implicitly ensures that the feature embeddings are invariant to the intervention, i.e. data augmentation. However, feature collapse is possible if a naïve implementation is used as can be seen in our evaluation (Table 1), where the naïve implementation of contrastive loss (ReCoRe-D) completely fails ($<$1% success rate) for model-based RL. Mitrovic et al. [24] proposed a regularizer based on KL-divergence that matches the distributions among the augmentations, which stabilizes contrastive learning under model-free RL settings and slightly improved the performance over CURL [20]. On the contrary, we propose a regularizer in the form of an auxiliary task to explicitly enforce the invariant feature learning. For example in the navigation task, we utilize depth predictions to extract the geometric features needed for navigation as they do not depend on textures as shown in Figure 1. We emphasize that depth is only required for training but not for deployment since it works as a regulariser. Furthermore, in cases where depth is not available other auxiliary tasks such as image denoising, segmentation or optical flow prediction can be utilized for regularization, enabling a wider applicability of the proposed model. Importantly, our setup allows us to employ contrastive learning in model-based RL settings, which improves the sample efficiency and helps with Out-of-Distribution (OoD) generalization.
@@
In summary, we propose a Regularized Contrastive Representation learning (ReCoRe) approach to the world model. The proposed variant of the world model can extract and predict robust invariant features (Figure 2). ReCoRe is verified on the point goal navigation task from Gibson [37] and iGibson 1.0 [30] as well as on the DeepMind Control suite (DMControl) [33]. Thus, our main contributions are:
@@
    We show that contrastive unsupervised representation learning can significantly improve OoD generalization of world model based reinforcement learning (Table 1, ReCoRe vs. DreamerV2).
@@
    We propose an intervention-invariant regularizer that learns the invariant features (Section 3.1.1) and is shown to be crucial in preventing feature collapse of contrastive learning (Table 1, ReCoRe vs. ReCoRe-D).
@@
    Through extensive experiments, we showcase that our approach outperforms state-of-the-art RL models (including language guided foundation model Grounding DINO [23]) on out-of-distribution generalization (Table 1) and sim-to-real transfer of learned features (Table 2). We further show that even for in-distribution evaluation our approach outperforms model-free reinforcement learning approaches (Table 3) which is difficult for other model-based learning approaches.
@@ 2 Related Work
Unsupervised Representation Learning. Learning reusable feature representations from large unlabeled data is a fundamental challenge in machine learning. In the context of computer vision, one can leverage unlabeled images and videos to learn good intermediate representations, which can be useful for a wide variety of downstream tasks. Recently, VAE [18] has been a preferred approach for representation learning in model-based RL [10]. Since VAE does not make any additional consideration of downstream tasks, invariant representation learning with contrastive loss has shown more promising results [1, 20]. Self-supervised learning formulates representation learning as a supervised loss function between different transformations of data. In image-based learning self-supervision can be formulated using different image augmentations, for example, image distortion and rotation [7, 3]. We also use different data augmentation techniques to learn the invariant features using contrastive loss. Recently, transformer-based visual models such as DINO [2] have been shown to intrinsically capture robust object representations through self-supervision. Combining such models with language models pretrained on a large corpus of data [22, 23] have resulted in powerful representations that generalize well to diverse environments. However, inference on such large models is computationally expensive which makes it difficult to deploy on low-budget robots. We also compare our proposed approach to Grounding DINO [23] features, which we believe is the strongest baseline, and showcase superior results.
Contrastive Learning. Representation learning methods based on contrastive loss [4] have achieved state-of-the-art performance on face verification tasks. These methods use a contrastive loss to learn representations invariant to data augmentation [3, 16]. Given a list of input samples, contrastive loss forces samples from the same class to have similar embeddings and different ones for different classes. Since class labels are not available in the unsupervised setting, contrastive loss forces similar embedding for the augmented version of the same sample and different ones for different samples. There are several ways of formulating the contrastive loss such as Siamese [4], InfoNCE [34], and SimCLR [3]. In this work, we chose InfoNCE [34] for our contrastive loss
@@
Model-based RL. The human brain discovers the underlying hidden causes of an observation. Those internal representations of the world influence how agents infer which actions will lead to a higher reward [15]. An early example of this idea was put forward by Sutton [32], where future hallucination samples rolled out from the learned world model are used in addition to the agent’s interactions for sample efficient learning. Further, planning through the world model has been successfully demonstrated in the world model by Ha et al. [10] and DreamerV2 by Hafner et al. [13]. Recently, replacing state extraction [29] and dynamic prediction [27] using transformer architecture is a popular direction, which further improves the results [29]. Masked World Models (MWM) combines the transformer-based masked autoencoder with DreamerV2. Other Dreamer variants DayDreamer [36] and DreamerV3 [14] aim to scale up DreamerV2 architecture for physical robots and Atari games. Specifically, DayDreamer [36] trains DreamerV2 on physical robots, while DreamerV3 [14] proposes design choices (e.g. symlog scaling and EMA regularization) that help in scaling up DreamerV2 to several domains. In our work, we propose to learn invariant features to improve OoD generalization and sample efficiency further. While we utilize a similar architecture as DreamerV2 for policy and world model, our method is complementary to DreamerV3 and can take advantage of their design choices.
Sample Efficiency. Joint learning of auxiliary tasks with model-free RL makes them competitive with model-based RL in terms of sample efficiency. For example, the recently proposed model-free RL method called CURL [20] added contrastive loss as an auxiliary task and outperformed the state-of-the-art model-based RL method called Dreamer [12]. Also, two recent works using data augmentation for RL called RAD [19] and DrQ [38] outperform CURL without using an auxiliary contrastive loss. These results warrant that if an agent has access to a rich stream of data from the environments, an additional regularizer is unnecessary since directly optimizing the policy objective is better than optimizing multiple objectives. However, we do not have access to a rich stream of data for many complex problems, hence sample efficiency still matters. Further, these papers do not consider the effect of regularizers in the form of auxiliary tasks and unsupervised representation learning for model-based RL, which is the main focus of our work.
@@ 3 ReCoRe-based World Model
We consider the visual control task as a finite-horizon partially observable Markov decision process (POMDP). We denote observation space, action space and time horizon as $\mathcal{O}$, $\mathcal{A}$ and $\mathcal{T}$ respectively. An agent performs continuous actions $a_{t}\sim p(a_{t}|o_{\leq t},a_{<t})$, and receives observations and scalar rewards $o_{t},r_{t}\sim p(o_{t},r_{t}|o_{<t},a_{<t})$ from the unknown environment. The goal of an agent is to maximize the expected total rewards $E_{p}(\sum_{t=1}^{T}r_{t})$. In the following sections, we detail our proposed model.
@@ 3.1 World Model Design
We propose our Regularized Contrastive Representation learning (ReCoRe) technique to learn the world model. The data flow diagram of our proposed world model with explicitly regularized invariant feature learning is shown in Figure 2. Our method consists of four main components: (i) invariant representation learning module, (ii) intervention-invariant regularizer, (iii) latent dynamics model, and (iv) the controller. Next, we describe the components in detail.
@@ 3.1.1 Invariant Representations Learning Module
Extracting representations that are invariant to appearance variations from image observations is a key component of our model. This is crucial for improving robustness and out-of-distribution generalization of the model in the real world. We learn these invariant representations by maximizing agreement between different style interventions of the same observation via a contrastive loss in the latent feature space. The world model optimizes feature learning and controller separately to improve the sample efficiency and simplify controller learning. Motivated by the fact that most of the complexity of model-based RL approaches resides in the world model (i.e. the feature extraction and the dynamics model), we hypothesize that an additional supervisory signal from an auxiliary task helps to learn a better state representation. In this work, we used InfoNCE [34] style loss to learn invariant features. Hence, our encoder takes RGB observations ($o_{t}$) as inputs and decodes rewards ($r_{t}$) as well as depth ($o_{t}^{d}$) as auxiliary outputs during the training phase. The invariant feature learning is enforced by contrastive loss ($\mathcal{L}^{q}_{t}$). The proposed invariant features learning technique has the following three sub-modules:
  • A style intervention module that utilizes data augmentation. We use spatial jitter, Gaussian blur, color jitter, grayscale and cutout data augmentation techniques for style intervention. Spatial jitter is implemented by first padding and then performing random crop. Given any observation $o_{t}$, our style intervention module randomly transforms it into two correlated views of the same observations, used for contrastive learning. All the hyperparameters are provided in the appendix.
  • We use an encoder network that extracts representations from augmented observations, which is $\tilde{s_{t}}=encoder(o_{t})$. The encoder is optimized using contrastive loss (Equation 1), which by construction makes the encoder invariant to the augmentations. We additionally employ an EMA regularization [16] on the encoder for stabilizing training. The hyperparameters of the encoder are provided in the appendix.
  • Contrastive loss is defined for a contrastive prediction task, which can be explained as a differentiable dictionary lookup task. Given a query observation $q$ and a set of keys $K=\{k_{0},k_{1},...k_{2B}\}$ of length $2B$, with known positive $\{k_{+}\}$ and negative $\{k_{-}\}$ keys, the aim of contrastive loss is to learn a representation in which positive sample pairs stay close to each other while negative ones are far apart. In contrastive learning $q$, $K$, $k_{+}$ and ${k_{-}}$ are also known as anchors, targets, positive and negative samples. We use bilinear products for projection head $W$ and InfoNCE loss for contrastive learning [34], which enforces the desired similarity in the embedding space: